Prediction for M_{W} vs M_{top}

Can use all LEP data and low-energy data

• Have to assume MHiggs

• Clearly sensitive test! – point represents CDF accuracy after Run II

Effect of Underlying Event

Have to understand effect of underlying event on M_{T}

• Enters in two ways:

1. Modifies u⃗ (and therefore E̸T)

2. Modifies electron energy measurement

Can study u⃗ effects by looking at Z events

• Z momentum well measured by dilepton final state

• Can determine calorimeter response to given recoil

How Well Can We Do?

The statistical errors will decrease ×2

• Already have ×5 more data

• Will come done to < ∼ 100 MeV/channel

Systematic uncertainties are the challenge

• Experience has shown them to also scale with sample size – All determined by other data

• Should hope to get to < ∼ 50 MeV/c2